OpenStudy (anonymous):

A carpenter is building a rectangular room with a fixed perimeter of 64ft. What dimensions would yield the maximum area? What is the maximum area?

OpenStudy (anonymous):

let x and y be the dimensions of room. 2x+2y=64 y=32-x area A=x*y=x(32-x)=32x-x^2 find dA/dx and find critical values.